[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Jun 26, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue Jun 27 02:57:08 UTC 2017


HSU Shift Log for Mon Jun 26 2017 22:57:01 GMT-0400 (EDT)

Shift: R

Forecaster: D Brown

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues: Interactive cone graphic did not update properly  
for Dora advisory 8 (21z). Dave Z. was contacted and fixed the issue.  
Should work for subsequent advisories.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170627/22aca70a/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list